Is this statement true or false ? provide reason for your answer 


d/dx[ x2∫(π/2) e2cost dt ] = -2xe2(cosx^2)


1
Expert's answer
2022-05-30T15:12:39-0400
"\displaystyle\int_{x}^{2}\dfrac{\pi}{2}e^{2\cos t}dt=-\displaystyle\int_{2}^{x}\dfrac{\pi}{2}e^{2\cos t}dt"

"\dfrac{d}{dx}(\displaystyle\int_{x}^{2}\dfrac{\pi}{2}e^{2\cos t}dt)=-\dfrac{d}{dx}(\displaystyle\int_{2}^{x}\dfrac{\pi}{2}e^{2\cos t}dt)"

"=-\dfrac{\pi}{2}e^{2\cos x}"

The statement is false.